You are given the following data in billions of dollars on your country’s economy:

Investment – 140, Government spending = 120, Exports = 80, Imports = 60, Autonomous consumption = 100, marginal propensity to save = 0.30.

a) Compute the consumption and saving functions for this economy and indicate the assumptions used. (4 Marks)

b) Compute the equilibrium level of income. (4 Marks)

c) Determine the equilibrium level of consumption and savings. (2 Marks)

d) Define and compute the multiplier. (2 Marks)

e) You are given that the full employment level of national income is 733 billion. Would this economy be experiencing deflationary of inflationary gap? Why? What is the magnitude of the gap and what could the government do to remove it? (4 Marks)

f) The government decided to increase its spending by 50 billion. Compute the equilibrium level of income and explain why income has or has not changed from the original level. (4 Marks)
1
Expert's answer
2018-11-09T15:49:09-0500

I = 140, G = 120, X = 80, M = 60, Ca = 100, s = 0.30.

a) Consumption function is C = Ca + c*(Y - T) and saving function for this economy is S = I. The assumptions used are S = I, T = G, c + s = 1.

b) The equilibrium level of income is:

Y = C + I + G + (X - M) = 100 + (1 - 0.3)*(Y - 120) + 140 + 120 + (80 - 60),

Y = 0.7Y + 296,

Y = 986.67.

c) The equilibrium level of consumption is:

C = 100 + 0.7*(986.67 - 120) = 706.67, saving is S = I = 140.

d) The multiplier is m = 1/s = 1/0.3 = 3.33.

e) If the full employment level of national income is 733 billion, then Y > Yp, so this economy is experiencing inflationary gap. The government can increase taxes or decrease government spending to remove it.

f) If the government decided to increase its spending by 50 billion, then the equilibrium level of income will increase by 50*3.33 = 166.67 million due to multiplier effect.